# Regional chemical potential analysis for material surfaces

### v1

@@ VI.4 C(001) diamond surface
The dangling bonds and double bonds in molecules are usually analyzed using other properties such as the partial density of states, the electron charge density [46], the frontier orbitals (Fukui function) [47], etc. However, these properties are sometimes not sufficient to analyze the surface of the materials. Figures 7 (a) and (b) show a structure of the C dimer ribbon on the C(001) diamond surface, which was recently investigated in great detail by sophisticated AFM experiments [20, 19]. The AFM images observed by the experiments revealed that the force between the AFM tip and the dimers of the ends of the C dimer ribbon is stronger than the force between the tip and the inner dimers [20]. Figures 7 (c)-(e) and Fig. 8 (a) show the partial density of states and electron charge density change $\Delta\rho(\vec{r})$, which is approximately proportional to the Fukui function $f(\vec{r})\equiv\frac{\partial\rho(\vec{r})}{\partial N}\propto\Delta\rho(\vec{r})$. There is little difference among C dimer atoms from the point of view of the partial density of states (PDOS) and the approximated Fukui function, so that they cannot give a clear picture of the electron-donating ability.
On the other hand, the distribution of $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ as shown in Fig. 8 (b) reveals that each C atom of the end dimers forms a double bond with a C atom of the substrate below. This is consistent with the discussion of the bond length [20]. The bond lengths of the corresponding double bond are approximately 1.38 - 1.40 Å, while those of the other bonds in the dimer ribbon are approximately 1.50 Å as shown in Figs. 7 (a) and (b). The red region shown in Fig. 8 (b) corresponds to the high local RCP $\mu_{R}^{\tau}(\vec{r})$ originating from the C-C double bonds, which are highly electron-donating regions. Electrons are more likely to be supplied to outer regions in their electron-donating regions, which can form covalent bonding with an approaching tip of atomic force spectroscopy. As a result, higher-resolution AFM images are observed at the regions. This result is consistent with the topology image of near-contact AFM experiments using an active Si probe [20, 19].
@@
Figure 7: The top view (a) and side view (b) of a geometrical structure of the 5 carbon dimer ribbon on C(001) diamond surface. Colors of the bonds represent the bond length. (c)-(e) The partial DOSs of three C atoms of the dimer ribbon. The area colored in orange represents the energy window for RCP $\mu_{R}^{\tau}(\vec{r})$ in Fig. 8
Figure 8: (a) The distribution of electron density of the 5 carbon dimer ribbon on C(001) diamond surface in the energy range between -2 eV and 0 eV (approximated Fukui function). (b) The corresponding RCP $\mu_{R}^{\tau}(\vec{r})$ on the surface. The drawing area is cropped from the actual system.
@@ VII Conclusions
The local RCP at the material surface $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ is a measure of how much electron transfer from the surface. It has been revealed that there are two types of electron transfers: “electron transfer from the material surface to the tip” (ionic bonding picture) and “electron transfer from the material surface or tip to the bonding region” (covalent bonding picture). Both cases of electron transfers are related with molecular absorption and interactions between a tip and a surface. In particular, we have focused on covalent bonding and discussed the relationship between the chemical bonding force and $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ using simple model calculations. Furthermore, DFT calculations of several molecules and C(001) diamond surfaces have been performed to visualize spatial distribution of the RCPs $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$. As a result, we have successfully visualized dangling bonds and double bonds, which are regions with high electron-donating ability. In other words, the RCPs $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ have visualized electron-donating ability at the surface where covalent bonds tend to be formed potentially. These results indicate that the RCP $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ has potential to analyze high-resolution AFM images, which are observed by chemical bonding forces. Furthermore, incorporating atomic height information and surface deformations that occur during AFM measurements remains an important subject for future work.
@@ VIII Acknowledgement
The computation in this work has been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo (ISSPkyodo-SC-2023-Ba-0073,2024-Ba-0053).
@@ Appendix A Approximation for DFT
It is not easy to calculate the electronic physical quantities using a quantum field theoretical state vector. Instead we approximate the physical quantities defined in the framework of the quantum field theory using density matrix obtained by two component relativistic DFT.
@@
The DFT approximation for the density matrix gives a approximated expectation value of electron number density.
@@
The expectation value of the electron number is given as
$n(\vec{r})$ $\equiv\langle\hat{n}(\vec{r})\rangle$ (42)
$\approx\frac{1}{V_{B}^{2}}\int_{B}d^{3}\vec{k}\int_{B}d^{3}\vec{k}^{\prime}\sum_{i,i^{\prime}=0}^{\infty}{\psi}_{i\vec{k}}^{\dagger}(\vec{r}){\psi}_{i^{\prime}\vec{k}^{\prime}}(\vec{r})\left\langle\hat{e}_{i\vec{k}}^{\dagger}\hat{e}_{i^{\prime}\vec{k}^{\prime}}\right\rangle_{\rm DFT},$ (43)
@@
${\tau}^{Slm}_{ei\vec{k}}(\vec{r})$ $\approx\frac{\hbar^{2}}{4m_{e}}\left[{\psi}_{i\vec{k}}^{\dagger}D_{el}D_{em}{\psi}_{i\vec{k}}-\left(D_{el}{\psi}_{i\vec{k}}\right)^{\dagger}D_{em}{\psi}_{i\vec{k}}\right]+h.c..$ (49)
@@
S1.1 Electronic transfer for ionic bonding
@@
S1.2 Energy decomposition and virial theorem
Total energy of the Rigged QED system (electron, photon, and nuclei) is defined as the integral of the 0-0 component of the energy momentum tensor $T^{\mu\nu}$, which is consist of three contributions from electron field, nucleus field, and photon field.
@@
If the system is in an equilibrium stationary state, virial theorem can be applied and we obtain the following relation.
@@
$=\left\langle\int\hat{T}_{e\mu}{}^{\mu}d^{3}\vec{r}\right\rangle+\sum_{a}\left\langle\int\hat{T}_{a\mu}{}^{\mu}d^{3}\vec{r}\right\rangle$ (S7)
@@
$=\left\langle\int\left[m_{e}c^{2}\hat{\bar{\psi}}\hat{\psi}\right]d^{3}\vec{r}\right\rangle.$ (S10)
The four-component $\psi$ can be expressed by two-component $\psi^{L}$ and $\psi^{S}$. The $\psi^{S}$ can be expressed by $\psi^{L}$ under the time-independent stationary state as follows.
$\hat{\psi}=\begin{pmatrix}\hat{\psi}^{L}\\ \hat{\psi}^{S}\end{pmatrix},\ \ \ \hat{\psi}^{S}\approx\frac{-i\hbar K}{2m_{e}c}\sigma^{m}\partial_{m}\hat{\psi}^{L}.$ (S11)
The above relation is obtained by the Dirac equation. Under the non-relativistic limit, $K\approx 1$. For simplicity, we ignored the vector potential.
@@
$E_{e}^{NR}$ $=E_{e\rm Rigged\ QED}-\left\langle\int m_{e}c^{2}\hat{\psi}^{\dagger}\hat{\psi}d^{3}\vec{r}\right\rangle,$
@@
$\approx\left\langle\int\left[-2m_{e}c^{2}\left(\frac{\hbar^{2}K^{2}}{4m_{e}^{2}c^{2}}\left(\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\sigma^{l}\sigma^{m}\partial_{m}\hat{\psi}^{L}\right)\right]d^{3}\vec{r}\right\rangle,$
$=\left\langle\int\left[-2m_{e}c^{2}\left(\frac{\hbar^{2}K^{2}}{4m_{e}^{2}c^{2}}\left(\left(\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\partial_{l}\hat{\psi}^{L}+i\varepsilon_{lmn}\left(\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\sigma^{n}\partial_{m}\hat{\psi}^{L}\right)\right)\right]d^{3}\vec{r}\right\rangle,$
$=\left\langle\int\left[-2m_{e}c^{2}\left(\frac{\hbar^{2}K^{2}}{4m_{e}^{2}c^{2}}\left(\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\partial_{l}\hat{\psi}^{L}\right)\right]d^{3}\vec{r}\right\rangle,$
$=\left\langle\int\left[-2m_{e}c^{2}\left(-\frac{\hbar^{2}K^{2}}{8m_{e}^{2}c^{2}}\left(\hat{\psi}^{L\dagger}\partial_{l}\partial_{l}\hat{\psi}^{L}+\left(\partial_{l}\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\hat{\psi}^{L}\right)\right)\right]d^{3}\vec{r}\right\rangle,$
$=\left\langle\int\left[\frac{\hbar^{2}K^{2}}{4m_{e}}\left(\left(\hat{\psi}^{L\dagger}\partial_{l}\partial_{l}\hat{\psi}^{L}+\left(\partial_{l}\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\hat{\psi}^{L}\right)\right)\right]d^{3}\vec{r}\right\rangle,$
$=\left\langle\int\left[-\frac{1}{2}\left(-\frac{\hbar^{2}K^{2}}{2m_{e}}\hat{\psi}^{L\dagger}\partial_{l}\partial_{l}\hat{\psi}^{L}+h.c.\right)\right]d^{3}\vec{r}\right\rangle,$
@@
$\approx-\frac{\hbar^{2}K^{2}}{4m_{e}}\hat{\psi}^{L\dagger}\partial_{l}\partial_{l}\hat{\psi}^{L}+h.c..$ (S14)
Similarly, the energy of nuclei part is written by kinetic energy of nuclei too. Therefore,
@@
Furthermore, the kinetic energy of electron can be expanded by minus half of integral over the trace of the electronic stress tensor density. Therefore,
$E_{\rm Rigged\ QED}$ $\approx-\langle\hat{T}_{e}\rangle,$ (S17)
$=\frac{1}{2}\int d^{3}\vec{r}\left\langle\sum_{k=1}^{3}\hat{\tau}_{e}^{kk}(\vec{r})\right\rangle.$ (S18)
This expression gives an energy decomposition scheme using electronic stress tensor density. The energy density constructed from electronic stress tensor density is defined as follows.
$\hat{\varepsilon}_{\tau}(x)$ $\equiv\frac{1}{2}\sum_{k=1}^{3}\hat{\tau}^{Skk}(\vec{r}\,),$ (S19)
@@
$\hat{\varepsilon}_{\tau}(\vec{r}\,)$ $\approx\frac{\hbar^{2}}{8m}\sum_{k=1}^{3}\left[\hat{\psi}^{L\dagger}\partial_{k}\partial_{k}\hat{\psi}^{L\dagger}-\left(\partial_{k}\hat{\psi}^{L}\right)^{\dagger}\partial_{k}\hat{\psi}^{L}+h.c.\right],$
@@
At the material surface, where the kinetic energy density equals zero,
$\hat{\varepsilon}_{\tau}({\rm surface})$ $=-\frac{\hbar^{2}}{8m}\sum_{k=1}^{3}\left[\partial_{k}\partial_{k}\left(\hat{\psi}^{L\dagger}\hat{\psi}^{L}\right)\right],$
@@
In AIM theory, $\langle\nabla^{2}\hat{n}\rangle$ is used for a measure of the localization at the bonding region.
S1.3 Local mechanical picture of covalent bonding

### v2

@@ VI.4 C(001) diamond surface
The dangling bonds and double bonds in molecules are usually analyzed using other properties such as the partial density of states, the electron charge density [48], the frontier orbitals (Fukui function) [49]. However, these properties are sometimes not sufficient to analyze the surface of the materials. Figures 7 (a) and (b) show a structure of the C dimer ribbon on the C(001) diamond surface, which was recently investigated in great detail by sophisticated AFM experiments [21, 20]. The AFM images observed by the experiments revealed that the force between the AFM tip and the dimers of the ends of the C dimer ribbon is stronger than the force between the tip and the inner dimers [21]. Figures 7 (c)-(e) and Fig. 8 (a) show the partial density of states and electron charge density change $\Delta\rho(\vec{r})$, which is approximately proportional to the Fukui function $f(\vec{r})\equiv\frac{\partial\rho(\vec{r})}{\partial N}\propto\Delta\rho(\vec{r})$. There is little difference among C dimer atoms from the point of view of the partial density of states (PDOS) and the approximated Fukui function, so that they cannot give a clear picture of the electron-donating ability.
Figure 7: The top view (a) and side view (b) of a geometrical structure of the 5 carbon dimer ribbon on C(001) diamond surface. Colors of the bonds represent the bond length. (c)-(e) The partial DOSs of three C atoms of the dimer ribbon. The area colored in orange represents the energy window for RCP $\mu_{R}^{\tau}(\vec{r})$ in Fig. 8.
However, the distribution of $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ as shown in Fig. 8 (b) reveals that each C atom of the end dimers forms a double bond with a C atom of the substrate below. This is consistent with the discussion of the bond length [21]. The bond lengths of the corresponding double bond are approximately 1.38 - 1.40 Å, while those of the other bonds in the dimer ribbon are approximately 1.50 Å as shown in Figs. 7 (a) and (b). The red region shown in Fig. 8 (b) corresponds to the high local RCP $\mu_{R}^{\tau}(\vec{r})$ originating from the C-C double bonds, which are highly electron-donating regions. Electrons are more likely to be supplied to outer regions in their electron-donating regions, which can form covalent bonding with an approaching tip of atomic force spectroscopy. As a result, higher-resolution AFM images are observed at the regions. This result is consistent with the topology image of near-contact AFM experiments using an active Si probe [21, 20].
@@
Figure 8: (a) The isosurface of electron density of the 5 carbon dimer ribbon on C(001) diamond surface in the energy range between -2 eV and 0 eV (approximated Fukui function). The isosurface value is 0.01 a.u. (b) The corresponding RCP $\mu_{R}^{\tau}(\vec{r})$ on the surface. The drawing area is cropped from the actual system.
@@ VII Conclusions
The local RCP at the electronic interface $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ is a measure of how much electron transfer from the surface. Two types of electron transfers have been discussed here: “electron transfer from the material surface to the tip” (ionic bonding picture) and “electron transfer from the material surface or tip to the bonding region” (covalent bonding picture). Both cases of electron transfers are related to molecular adsorption and interactions between a tip and a surface. In particular, we have focused on covalent bonding and discussed the relationship between the chemical bonding force and $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ using simple model calculations. Furthermore, DFT calculations of several molecules and C(001) diamond surfaces have been performed to visualize spatial distribution of the RCPs $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$. As a result, we have successfully visualized dangling bonds and double bonds, which are regions with high electron-donating ability. In other words, the RCPs $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ have visualized electron-donating ability at the surface where covalent bonds tend to be formed potentially. These results indicate that the RCP $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ has potential to analyze high-resolution AFM images, which are observed by chemical bonding forces. Furthermore, incorporating atomic height information and surface deformations that occur during AFM measurements remains an important subject for future work.
@@ VIII Acknowledgement
The computation in this work has been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo (ISSPkyodo-SC-2023-Ba-0073,2024-Ba-0053). The work of M.S. was supported by Grants-in-Aid for Scientific Research (22K12060).
@@ Appendix A Approximation for DFT
It is not easy to calculate the electronic physical quantities using a quantum field theoretical state vector. Instead we approximate the physical quantities defined in the framework of the quantum field theory using density matrix obtained by two-component relativistic or non-relativistic DFT.
@@
The DFT approximation for the density matrix gives an approximated expectation value of the electron number density.
@@
where $\epsilon_{i}$ is the $i$-th eigenenergy. The expectation value of the electron number density is given as
$n(\vec{r})$ $\equiv\langle\hat{n}(\vec{r})\rangle\approx\langle\hat{n}(\vec{r})\rangle_{\rm DFT},$ (43)
$\left\langle\hat{n}(\vec{r})\right\rangle_{\rm DFT}$ $=\frac{1}{V_{B}^{2}}\int_{B}d^{3}\vec{k}\int_{B}d^{3}\vec{k}^{\prime}\sum_{i,i^{\prime}=0}^{\infty}{\psi}_{i\vec{k}}^{\dagger}(\vec{r}){\psi}_{i^{\prime}\vec{k}^{\prime}}(\vec{r})$
×$\quad\times\left\langle\hat{e}_{i\vec{k}}^{\dagger}\hat{e}_{i^{\prime}\vec{k}^{\prime}}\right\rangle_{\rm DFT},$ (44)
@@
${\tau}^{Slm}_{ei\vec{k}}$ $\approx\frac{\hbar^{2}}{4m_{e}}\left[{\psi}_{i\vec{k}}^{L\dagger}D_{el}D_{em}{\psi}^{L}_{i\vec{k}}-\left(D_{el}{\psi}^{L}_{i\vec{k}}\right)^{\dagger}D_{em}{\psi}^{L}_{i\vec{k}}\right]+h.c.$ (49)
For the linear combination of the atomic orbital (LCAO) method, the function $\psi^{L}_{i\vec{k}}(\vec{r})$ can be written by $\psi_{\sigma i}(\vec{k},\vec{r})$, which is expanded using expansion coefficients $c_{\sigma i,p\alpha}(\vec{k})$ and atomic orbitals $\phi_{p\alpha}(\vec{r})$ as follows.
$\psi^{L}_{i\vec{k}}(\vec{r})=\begin{pmatrix}\psi_{\uparrow i}(\vec{k},\vec{r})\\ \psi_{\downarrow i}(\vec{k},\vec{r})\end{pmatrix},$ (50)
where
$\psi_{\sigma i}(\vec{k},\vec{r})=\frac{1}{N}\sum_{n}^{N}e^{i\vec{T}_{n}\cdot\vec{k}}\sum_{p,\alpha}c_{\sigma i,p\alpha}(\vec{k})\,\phi_{p\alpha}(\vec{r}-\vec{R}_{p}-\vec{T}_{n}).$ (51)
Here, $\vec{T}_{n}$ is a lattice vector, $N$ is the number of a cell, $p(q)$ is a site index corresponding to a position $\vec{R}_{p}(\vec{R}_{q})$, $\sigma$ (either $\uparrow$ or $\downarrow$) denotes the spin index, and $\alpha$ and $\beta$ are the orbital indices. Then, Eq. (14) ignoring the vector potential is expanded as
${\tau}^{Slm}_{ei\vec{k}}(\vec{r})$ $\approx\frac{\hbar^{2}}{4m_{e}N}\sum_{n}^{N}\sum_{\sigma}\rho_{\sigma\sigma,p\alpha q\beta}^{(\vec{T}_{n})}(\vec{k})\sum_{j}^{2}\omega^{(j)lm(\vec{T}_{n})}_{p\alpha q\beta}(\vec{r})+h.c.,$ (52)
where
$\rho_{\sigma\sigma^{\prime},p\alpha q\beta}^{(\vec{T}_{n})}(\vec{k})$ $\equiv e^{i\vec{T}_{n}\cdot\vec{k}}c_{\sigma i,p\alpha}^{*}(\vec{k})c_{\sigma^{\prime}i,q\beta}(\vec{k}),$ (53)
$\omega^{(1)lm(\vec{T}_{n})}_{p\alpha q\beta}(\vec{r})$ $\equiv\phi^{*}_{p\alpha}\left(\vec{r}-\vec{R}_{p}\right)\partial_{l}\partial_{m}\phi_{q\beta}\left(\vec{r}-\vec{R}_{q}-\vec{T}_{n}\right),$ (54)
$\omega^{(2)lm(\vec{T}_{n})}_{p\alpha q\beta}(\vec{r})$ $\equiv-\partial_{l}\phi^{*}_{p\alpha}\left(\vec{r}-\vec{R}_{p}\right)\partial_{m}\phi_{q\beta}\left(\vec{r}-\vec{R}_{q}-\vec{T}_{n}\right).$ (55)
For the plane wave (PW) method, $\psi_{\sigma i}(\vec{k},\vec{r})$ is expanded by the coefficient $C_{\sigma i,\vec{G}}(\vec{k})$ and the plane wave $e^{i\vec{k}\cdot\vec{r}}$ as
$\psi_{\sigma i}(\vec{k},\vec{r})=\frac{1}{\sqrt{\Omega}}\sum_{\vec{G}}C_{\sigma i,\vec{G}}(\vec{k})e^{i(\vec{k}+\vec{G})\cdot\vec{r}},$ (56)
where $\vec{G}$ is a reciprocal lattice vector, and $\Omega$ is a normalization factor. Then, Eq. (14) ignoring the vector potential is expanded as
${\tau}^{Slm}_{ei\vec{k}}(\vec{r})$ $\approx\frac{\hbar^{2}}{4m_{e}\sqrt{\Omega}}\sum_{\vec{G},\vec{G}^{\prime}}\sum_{\sigma}\rho_{\sigma\sigma}^{\vec{G},\vec{G^{\prime}}}(\vec{k})\omega^{lm}_{\vec{k},\vec{G},\vec{G}^{\prime}}(\vec{r})+h.c.,$ (57)
where
$\rho_{\sigma\sigma^{\prime}}^{\vec{G},\vec{G}^{\prime}}(\vec{k})$ $\equiv C_{\sigma i,\vec{G}}^{*}(\vec{k})C_{\sigma^{\prime}i,\vec{G}^{\prime}}(\vec{k}),$ (58)
$\omega^{(1)lm}_{\vec{k},\vec{G},\vec{G}^{\prime}}(\vec{r})$ $\equiv-(k^{l}+G^{\prime l})(k^{m}+G^{\prime m})e^{i(\vec{G}^{\prime}-\vec{G})\cdot\vec{r}},$ (59)
$\omega^{(2)lm}_{\vec{k},\vec{G},\vec{G}^{\prime}}(\vec{r})$ $\equiv-(k^{l}+G^{l})(k^{m}+G^{\prime m})e^{i(\vec{G}^{\prime}-\vec{G})\cdot\vec{r}},$ (60)
$\omega^{lm}_{\vec{k},\vec{G},\vec{G}^{\prime}}(\vec{r})$ $\equiv\omega^{(1)lm}_{\vec{k},\vec{G},\vec{G}^{\prime}}(\vec{r})+\omega^{(2)lm}_{\vec{k},\vec{G},\vec{G}^{\prime}}(\vec{r}),$ (61)
$=-(2k^{l}+G^{l}+G^{\prime l})(k^{m}+G^{\prime m})e^{i(\vec{G}^{\prime}-\vec{G})\cdot\vec{r}}.$ (62)
In this study, we used Eq. (52) to calculate the RCP based on the LCAO method by interpreting the atomic orbitals as local pseudo-atomic basis functions.
@@
S1 Electronic transfer for ionic bonding
@@
S2 Energy decomposition and virial theorem
Total energy of the Rigged QED system (electron, photon, and nuclei) is defined as the integral of the 0-0 component of the energy momentum tensor $T^{\mu\nu}$, which consists of three contributions from electron field, nucleus field, and photon field.
@@
If the system is in an equilibrium stationary state, the virial theorem can be applied, and we obtain the following relation.
@@
$=\left\langle\int\hat{T}_{e\mu}{}^{\mu}d^{3}\vec{r}\right\rangle+\sum_{a}\left\langle\int\hat{T}_{a\mu}{}^{\mu}d^{3}\vec{r}\right\rangle.$ (S7)
@@
$=\left\langle\int\left[m_{e}c^{2}\hat{\bar{\psi}}\hat{\psi}\right]d^{3}\vec{r}\right\rangle,$ (S10)
where $\bar{\psi}\equiv\hat{\psi^{\dagger}\gamma^{0}}$. The matrices $\gamma^{0}$ and $\gamma^{k}$ are the 4 × 4 gamma matrices. The four-component $\psi$ can be expressed by two-component $\psi^{L}$ and $\psi^{S}$. The $\psi^{S}$ can be expressed by $\psi^{L}$ in the time-independent stationary state as follows [16].
$\hat{\psi}=\begin{pmatrix}\hat{\psi}^{L}\\ \hat{\psi}^{S}\end{pmatrix},\ \ \ \hat{\psi}^{S}\approx\frac{-i\hbar\bar{K}}{2m_{e}c}\sigma^{m}\partial_{m}\hat{\psi}^{L},\ \ \ \bar{K}=\frac{2m_{e}c^{2}}{2m_{e}c^{2}+E_{e\rm Rigged\ QED}-Z_{e}eA_{0}},$ (S11)
where $A_{0}$ is the scalar potential. The above relation is obtained by the Dirac equation. Under the non-relativistic limit, $\bar{K}\approx 1$. For simplicity, we ignored the vector potential.
@@
$E_{e\rm Rigged\ QED}-\left\langle\int m_{e}c^{2}\hat{\psi}^{\dagger}\hat{\psi}d^{3}\vec{r}\right\rangle,$
@@
$\approx E_{e}^{NR},$ (S12)
$E_{e}^{NR}$ $=\left\langle\int\left[-2m_{e}c^{2}\left(\frac{\hbar^{2}\bar{K}^{2}}{4m_{e}^{2}c^{2}}\left(\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\sigma^{l}\sigma^{m}\partial_{m}\hat{\psi}^{L}\right)\right]d^{3}\vec{r}\right\rangle,$
$=\left\langle\int\left[-2m_{e}c^{2}\left(\frac{\hbar^{2}\bar{K}^{2}}{4m_{e}^{2}c^{2}}\left(\left(\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\partial_{l}\hat{\psi}^{L}+i\varepsilon_{lmn}\left(\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\sigma^{n}\partial_{m}\hat{\psi}^{L}\right)\right)\right]d^{3}\vec{r}\right\rangle,$
$=\left\langle\int\left[-2m_{e}c^{2}\left(\frac{\hbar^{2}\bar{K}^{2}}{4m_{e}^{2}c^{2}}\left(\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\partial_{l}\hat{\psi}^{L}\right)\right]d^{3}\vec{r}\right\rangle,$
$=\left\langle\int\left[-2m_{e}c^{2}\left(-\frac{\hbar^{2}\bar{K}^{2}}{8m_{e}^{2}c^{2}}\left(\hat{\psi}^{L\dagger}\partial_{l}\partial_{l}\hat{\psi}^{L}+\left(\partial_{l}\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\hat{\psi}^{L}\right)\right)\right]d^{3}\vec{r}\right\rangle,$
$=\left\langle\int\left[\frac{\hbar^{2}\bar{K}^{2}}{4m_{e}}\left(\left(\hat{\psi}^{L\dagger}\partial_{l}\partial_{l}\hat{\psi}^{L}+\left(\partial_{l}\partial_{l}\hat{\psi}^{L}\right)^{\dagger}\hat{\psi}^{L}\right)\right)\right]d^{3}\vec{r}\right\rangle,$
$=\left\langle\int\left[-\frac{1}{2}\left(-\frac{\hbar^{2}\bar{K}^{2}}{2m_{e}}\hat{\psi}^{L\dagger}\partial_{l}\partial_{l}\hat{\psi}^{L}+h.c.\right)\right]d^{3}\vec{r}\right\rangle,$
@@
$\approx-\frac{\hbar^{2}\bar{K}^{2}}{4m_{e}}\hat{\psi}^{L\dagger}\partial_{l}\partial_{l}\hat{\psi}^{L}+h.c.$ (S15)
Similarly, the energy of the nuclei part is written by the kinetic energy of the nuclei too. Therefore,
@@
Furthermore, the kinetic energy of the electron can be expanded by minus half of the integral over the trace of the electronic stress tensor density. Therefore,
$E_{\rm Rigged\ QED}$ $\approx-\langle\hat{T}_{e}\rangle=\frac{1}{2}\int d^{3}\vec{r}\left\langle\sum_{k=1}^{3}\hat{\tau}_{e}^{kk}(\vec{r})\right\rangle.$ (S18)
This expression gives an energy decomposition scheme using electronic stress tensor density. The regional energy density operator constructed from the electronic stress tensor density operator is defined as follows.
$\hat{\varepsilon}_{\tau}(\vec{r})$ $\equiv\frac{1}{2}\sum_{k=1}^{3}\hat{\tau}^{Skk}(\vec{r}\,),$ (S19)
@@
$\hat{\varepsilon}_{\tau}(\vec{r}\,)\approx\hat{\varepsilon}^{\rm NR}_{\tau}(\vec{r}\,)$ $=\frac{\hbar^{2}}{8m}\sum_{k=1}^{3}\left[\hat{\psi}^{L\dagger}\partial_{k}\partial_{k}\hat{\psi}^{L\dagger}-\left(\partial_{k}\hat{\psi}^{L}\right)^{\dagger}\partial_{k}\hat{\psi}^{L}+h.c.\right],$
@@
At the electronic interface, where the kinetic energy density equals zero,
$\hat{\varepsilon}^{\rm NR}_{\tau}({\rm surface})$ $=-\frac{\hbar^{2}}{8m}\sum_{k=1}^{3}\left[\partial_{k}\partial_{k}\left(\hat{\psi}^{L\dagger}\hat{\psi}^{L}\right)\right],$
@@
In AIM theory, $\langle\nabla^{2}\hat{n}\rangle$ is used as a measure of localization in the bonding region.
S3 Local mechanical picture of covalent bonding